# Repeating Flares, X-ray Outbursts and Delayed Infrared Emission: A Comprehensive Compilation of Optical Tidal Disruption Events

### v1

@@ 5.3 Similarities and specific cases in the repeating flare sample
From our main catalogue, 12 TDEs exhibit more than one flare. While half of these TDEs have been extensively studied in previous works, 3 of the remaining 6 sources (AT 2020acka, AT 2019teq, and AT 2019ehz) have appeared in past sample studies (van Velzen et al. 2021; Hammerstein et al. 2023; Yao et al. 2023) without their repeating nature being highlighted. Only recently were AT 2020acka’s and AT 2019ehz’s re-brightening features referenced by Guo et al. (2025) and Zhong (2025). Unfortunately, due to a lack of spectroscopic data, we cannot confirm the TDE nature of the flares in the latter half of our repeating flare sample. However, by comparing the rise times, decay times and durations of both flares we find that the timescales for AT 2022dbl, AT 2022agi, AT 2020vdq, AT 2019ehz, and AT 2019azh, are consistent within uncertainty with one another (see Appendix D). We are unable to compute the rise time of the second flare of AT 2021mhg and the decay time of the second flare of AT 2019teq due to substantial gaps in the respective light curves. Additionally, for AT2022exr and AT2021uvz, which are the double peak flares, as well as AT2020acka, we cannot evaluate the shape of their flares since they are overlapping. From a visual inspection, the rise time of the second flare of AT 2021mhg is clearly different and has been identified as a Type Ia SN. Moreover, the flares in AT 2020acka and AT 2019teq, possibly originating from accretion disc emission and an unknown transient respectively, also display dissimilar shapes. The similarity of the repeating flares has also been previously noted by Makrygianni et al. (2025) for AT 2022dbl and Hinkle et al. (2021) for AT 2019azh.
@@
If the additional flares arise from pTDEs, one would expect their shapes to be similar since they would result from the disruption of the same star. However, the observed diversity in flare shapes suggests that underlying mechanisms may be at play: some flares could indeed be repeating pTDEs, while others may result from distinct events such as double TDEs or unrelated transient phenomena. For instance, the 18-year interval between flares in AT 2024pvu makes it unlikely that both events originate from the partial disruption of the same star. In this context, the overall shape of a flare may be influenced primarily by the properties of the SMBH rather than by the characteristics of the disrupted star. Moreover, Liu et al. (2025) found in their simulations that a sun-like star would produce multiple flares of increasing luminosity, consistent with AT 2020vdq, whose second flare is more luminous than the first. However, this picture is not consistent with AT 2022dbl, where the opposite behaviour is observed.
@@
Furthermore, 8 out of the 12 sources in our repeating flare sample exhibit X-ray detections with the exception of AT 2024pvu, AT 2022dbl, AT 2021mhg, and AT 2020acka, for which we obtained only upper limits. We observe that in some TDEs (AT 2022exr,,AT 2020vdq, AT 2019ehz, AT 2019teq, and AT 2018fyk), the X-ray emission follows the optical flare with only a short delay. In contrast, others (AT 2021uvz, AT 2022agi, and AT 2019azh) display significantly longer delays. Notably, for AT 2022exr and AT 2020vdq, X-rays are detected only after the second flare, while for AT 2018fyk they occur nearly simultaneously with the optical outbursts.
5.4 IR TDE candidate samples
In addition to the confirmed TDEs presented in Sect. 4.4, several studies have identified samples of infrared-selected TDE candidates. For example, Masterson et al. (2024) identified 18 candidates in nearby galactic nuclei (within 200 Mpc) based on distinctive mid-infrared light curve features. They divided these into two subsamples—a set of six candidates where the variability might be linked to an underlying AGN, and a set of 12 more robust candidates (including WTP14adbjsh, initially discussed in Panagiotou et al. 2023). Similarly, Jiang et al. (2021) examined IR flares in SDSS galaxies, identifying 137 mid-infrared outbursts, of which two have been proposed as TDE candidates. TDECat provides a sample of optical TDEs with an IR counterpart which can be used for comparison with IR-selected TDE candidate samples to investigate if a potential connection exists.
5.5 Correlation between X-ray and Infrared emission?
@@
As shown in Sect. 4.4, most TDEs exhibiting IR flares also display X-ray detections. In fact, 14 out of the 18 IR-flaring TDEs have been detected in X-rays. Moreover, in the sample of infrared-selected TDEs from the eROSITA sky presented by Masterson et al. (2024), 3 of the 8 objects (all from their gold sample) were detected in the X-rays at multiple epochs. These findings suggest a potential underlying correlation between IR and X-ray emission.
@@
Notably, both confirmed TDEs associated with astrophysical neutrinos, AT 2019dsg (Stein et al. 2021a) and AT 2017gge (Li et al. 2024), are included in our IR flare sample. It has been proposed that neutrino production in these events could be linked to radio emission (Stein et al. 2021a), infrared emission (Yuan et al. 2024), or X-ray emission (Li et al. 2024). However, further study is needed to clarify the association between neutrinos and TDEs, as well as the particle acceleration mechanisms that might lead to neutrino production.
@@
We can estimate the statistical significance of the correlation between IR and X-ray emission in our sample of optical TDEs, by considering a binomial calculation. The chance that a TDE has X-ray emission can be calculated as the detection fraction of the X-rays in our optical TDE sample (i.e. p=45/134=0.336). We can then test the null hypothesis (i.e. that each of the IR-flaring TDEs has an independent chance p=0.336 of X-ray detection), by computing the probability of seeing X-ray detections by chance
Since , we reject the null hypothesis, meaning that the observed correlation of X-ray and IR emission in our IR-emitting optical TDE sample is likely not by chance. We can also convert the probability to sigma levels for a one-tailed normal distribution, which gives a 3.6 significance.
5.6 X-ray TDE candidate samples
Previous works have studied the X-ray TDE candidate population (e.g. Auchettl et al. 2017; Sazonov et al. 2021; Khorunzhev et al. 2022). However, there is still uncertainty on the properties that distinguish X-ray TDEs from active galactic nuclei (AGN) activity. The TDEs in our X-ray sample, which also have an optical counterpart (except from the 3 jetted TDEs), provide a valuable opportunity to refine the classification criteria for X-ray TDEs that lack detectable optical emission. Moreover, the substantially larger number of X-ray bright optical TDEs now available enables more comprehensive comparative studies with the AGN population (e.g. Auchettl et al. 2018), which may ultimately help elucidate the intrinsic differences between these phenomena.
@@ 6 Conclusions
We have compiled a catalogue of 134 confirmed TDEs up until the end of 2024. We collected multi-wavelength photometry (X-ray, UV, optical, and infrared) along with publicly available spectra. The complete dataset is accessible via a dedicated GitHub repository and a Python application. Additionally, we created a list of strong TDE candidates.
@@
By implementing a custom Bayesian block algorithm, we estimated the duration, rise time, decay time, and their ratio for the optical flares of 101 TDEs in our sample. A total of 103 flares were analysed, including all flares from repeating pTDEs (AT 2022dbl and AT 2020vdq). We find that the distributions of the ratio, durations, rise times, and decay times are well described by a log-normal distribution.
@@
The best-fit line for vs is given by , where and , with 95% confidence intervals of 0.8374 to 1.1227 and -0.6859 to -0.0168 for and respectively. This result indicates that the majority of the TDE population is confined within the scatter of this best-fit line. We also find a positive correlation between and . These results, in combination with the log-normal distributions of the timescales, can be used to create artificial light curves for future statistical studies.
@@
Our spectral analysis reveals that the majority of TDEs belong to the TDE-H+He class, followed by the TDE-H class, with the TDE-He and TDE-featureless occurring at slightly lower frequencies. This distribution is consistent with expectations based on the predominance of main sequence stars in galactic centres. Although we computed timescale statistics for the different spectral classes, we did not identify any definite trends.
@@
We examined whether specific spectral classes tend to exhibit repeating flares, infrared emission, or X-ray outbursts. Our analysis reveals that the spectral class distribution within these TDE subsamples is unlikely to have arisen by random chance.
@@
We identified 4 new TDEs that show secondary flares in their optical light curves, namely AT 2024pvu, AT 2022exr, AT 2021uvz, and AT 2019teq. We further studied recently referenced TDEs displaying re-brightenings: namely AT 2020acka, and AT 2019ehz. We also compared the shape of the multiple flares for each TDE and found that their shapes are generally similar, excluding AT 2021mhg, AT 2020acka and AT 2019teq, which exhibit distinct differences.
@@
We observe a potential correlation between infrared and X-ray emission; most TDEs (14 out of 18) displaying infrared flares also exhibit X-ray detections. We plan to explore this finding further in a future study. Moreover, all reported coronal line emitting TDEs were found to exhibit both X-ray and IR emission.
This comprehensive (mostly optical) TDE catalogue not only provides a robust dataset for statistical studies and machine learning applications, but also paves the way for future population studies aimed at understanding the characteristics of still poorly understood TDEs. As we enter into the data ”rich” era, especially with forthcoming surveys like the LSST, which are expected to capture tens of TDEs each night, this work serves as an essential tool for preparing for the challenges and opportunities ahead.
@@ Acknowledgements.
The authors thank the anonymous referee for comments and suggestions that helped improve this work. D.A.L., A.P., and I.L. were funded by the European Union ERC-2022-STG - BOOTES - 101076343. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. N.G. acknowledge support by a grant from the Simons Foundation (00001470, NG). J.D., D.M., and Z.T. took part in this research under the auspices of the Science Internship Program at the University of California Santa Cruz. N.G. is grateful to Caitlyn Nojiri and Hannah Dykaar for their help. BHTOM has been based on the open-source TOM Toolkit by LCO and has been developed with funding from the European Union’s Horizon 2020 and Horizon Europe research and innovation programmes under grant agreements No 101004719 (OPTICON-RadioNet Pilot, ORP) and No 101131928 (ACME). We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts). This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The CSS survey is funded by the National Aeronautics and Space Administration under Grant No. NNG05GF22G issued through the Science Mission Directorate Near-Earth Objects Observations Program. The CRTS survey is supported by the U.S. National Science Foundation under grants AST-0909182. This publication makes use of data products from NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. This work utilises Matplotlib (Hunter 2007), NumPy (Harris et al. 2020) and SciPy (Virtanen et al. 2020).References
- Ajay et al. (2022) Ajay, Y., Pasham, D., & Guolo, M. 2022, The Astronomer’s Telegram, 15724, 1
- Alexander et al. (2016) Alexander, K. D., Berger, E., Guillochon, J., Zauderer, B. A., & Williams, P. K. G. 2016, ApJ, 819, L25
- Alexander et al. (2021) Alexander, K. D., Velzen, S. V., Miller-Jones, J., et al. 2021, Transient Name Server AstroNote, 24, 1
- Andreoni et al. (2022) Andreoni, I., Coughlin, M. W., Perley, D. A., et al. 2022, Nature, 612, 430
- Arcavi et al. (2014) Arcavi, I., Gal-Yam, A., Sullivan, M., et al. 2014, ApJ, 793, 38
- Auchettl et al. (2017) Auchettl, K., Guillochon, J., & Ramirez-Ruiz, E. 2017, ApJ, 838, 149
- Auchettl et al. (2018) Auchettl, K., Ramirez-Ruiz, E., & Guillochon, J. 2018, ApJ, 852, 37
- Bade et al. (1996) Bade, N., Komossa, S., & Dahlem, M. 1996, A&A, 309, L35
- Bellm et al. (2019) Bellm, E. C., Kulkarni, S. R., Graham, M. J., et al. 2019, PASP, 131, 018002
- Blagorodnova et al. (2019) Blagorodnova, N., Cenko, S. B., Kulkarni, S. R., et al. 2019, ApJ, 873, 92
- Blanton & Roweis (2007) Blanton, M. R. & Roweis, S. 2007, AJ, 133, 734
- Bloom et al. (2011) Bloom, J. S., Giannios, D., Metzger, B. D., et al. 2011, Science, 333, 203
- Brimacombe et al. (2015) Brimacombe, J., Brown, J. S., Holoien, T. W. S., et al. 2015, The Astronomer’s Telegram, 7910, 1
- Brown et al. (2015) Brown, G. C., Levan, A. J., Stanway, E. R., et al. 2015, MNRAS, 452, 4297
- Burrows et al. (2005) Burrows, D. N., Hill, J. E., Nousek, J. A., et al. 2005, Space Sci. Rev., 120, 165
- Burrows et al. (2011) Burrows, D. N., Kennea, J. A., Ghisellini, G., et al. 2011, Nature, 476, 421
- Capalbi et al. (2005) Capalbi, M., Perri, M., Saija, B., Tamburelli, F., & Angelini, L. 2005, Version, 1, 28
- Cendes et al. (2022) Cendes, Y., Alexander, K., Berger, E., et al. 2022, The Astronomer’s Telegram, 15728, 1
- Cenko et al. (2016) Cenko, S. B., Cucchiara, A., Roth, N., et al. 2016, ApJ, 818, L32
- Cenko et al. (2012) Cenko, S. B., Krimm, H. A., Horesh, A., et al. 2012, ApJ, 753, 77
- Chambers et al. (2016) Chambers, K. C., Magnier, E. A., Metcalfe, N., et al. 2016, arXiv e-prints, arXiv:1612.05560
- Charalampopoulos et al. (2022) Charalampopoulos, P., Leloudas, G., Malesani, D. B., et al. 2022, A&A, 659, A34
- Charalampopoulos et al. (2023) Charalampopoulos, P., Leloudas, G., Pursiainen, M., & Kotak, R. 2023, Transient Name Server AstroNote, 115, 1
- D’Elia et al. (2013) D’Elia, V., Perri, M., Puccetti, S., et al. 2013, A&A, 551, A142
- Drake et al. (2009) Drake, A. J., Djorgovski, S. G., Mahabal, A., et al. 2009, ApJ, 696, 870
- Fitzpatrick & Massa (2007) Fitzpatrick, E. L. & Massa, D. 2007, ApJ, 663, 320
- Flewelling et al. (2020) Flewelling, H. A., Magnier, E. A., Chambers, K. C., et al. 2020, ApJS, 251, 7
@@
- Freeman et al. (2001) Freeman, P., Doe, S., & Siemiginowska, A. 2001, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4477, Astronomical Data Analysis, ed. J.-L. Starck & F. D. Murtagh, 76–87

### v2

@@ 5.3 Similarities and specific cases in the repeating flare sample
From our main catalogue, 11 TDEs exhibit more than one flare. While six of these TDEs have been extensively studied in previous works, two of the remaining five sources (AT 2020acka, and AT 2019ehz) have appeared in past sample studies (van Velzen et al. 2021; Hammerstein et al. 2023; Yao et al. 2023) without their repeating nature being highlighted. Only recently were AT 2020acka’s and AT 2019ehz’s re-brightening features referenced by Guo et al. (2025) and Zhong (2025). Unfortunately, due to a lack of spectroscopic data, we cannot confirm the TDE nature of the flares in the latter half of our repeating flare sample. However, by comparing the rise times, decay times and durations of both flares we find that the timescales for AT 2022dbl, AT 2022agi, AT 2020vdq, AT 2019ehz, and AT 2019azh, are consistent within uncertainty with one another (see Appendix D). We are unable to compute the rise time of the second flare of AT 2021mhg due to substantial gaps in the respective light curve. Additionally, for AT 2022exr and AT 2021uvz, which are the double peak flares, as well as AT 2020acka, we cannot evaluate the shape of their flares since they are overlapping. From a visual inspection, the rise time of the second flare of AT 2021mhg is clearly different and has been identified as a Type Ia SN. Moreover, the flares in AT 2020acka, possibly originating from accretion disc emission, also displays dissimilar shapes. The similarity of the repeating flares has also been previously noted by Makrygianni et al. (2025) for AT 2022dbl and Hinkle et al. (2021) for AT 2019azh.
@@
If the additional flares arise from pTDEs, one could expect their shapes to be similar since they would result from the disruption of the same star (simulation papers that have modelled repeating pTDEs have yet to constrain rise times and decay times; e.g. Liu et al. 2025). However, the observed diversity in flare shapes suggests that underlying mechanisms may be at play: some flares could indeed be repeating pTDEs, while others may result from distinct events such as double TDEs or unrelated transient phenomena. Moreover, Liu et al. (2025) found in their simulations that a sun-like star would produce multiple flares of increasing luminosity, consistent with AT 2020vdq, whose second flare is more luminous than the first. However, this picture is not consistent with AT 2022dbl (e.g. Makrygianni et al. 2025), where the opposite behaviour is observed.
@@
Furthermore, seven out of the 11 sources in our repeating flare sample exhibit X-ray detections with the exception of AT 2024pvu, AT 2022dbl, AT 2021mhg, and AT 2020acka, for which we obtained only upper limits. We observe that in some TDEs (AT 2022exr, AT 2020vdq, AT 2019ehz, and AT 2018fyk), the X-ray emission follows the optical flare with only a short delay. In contrast, others (AT 2021uvz, AT 2022agi, and AT 2019azh) display significantly longer delays. Notably, for AT 2022exr and AT 2020vdq, X-rays are detected only after the second flare, while for AT 2018fyk they occur nearly simultaneously with the optical outbursts.
5.4 Possible correlation between X-ray and infrared emission
@@
As shown in Sect. 4.4, most TDEs exhibiting IR flares also display X-ray detections. In fact, 14 out of the 18 IR-flaring TDEs have been detected in X-rays. Moreover, in the sample of IR-selected TDEs from the eROSITA sky presented by Masterson et al. (2024), three of the eight objects (all from their gold sample) were detected in the X-rays at multiple epochs. These findings suggest a potential underlying correlation between IR and X-ray emission.
@@
Notably, both confirmed TDEs associated with astrophysical neutrinos, AT 2019dsg (Stein et al. 2021) and AT 2017gge (Li et al. 2024), are included in our IR flare sample. It has been proposed that neutrino production in these events could be linked to radio emission (Stein et al. 2021), IR emission (Yuan et al. 2024), or X-ray emission (Li et al. 2024). However, further study is needed to clarify the association between neutrinos and TDEs, as well as the particle acceleration mechanisms that might lead to neutrino production.
@@
The statistical significance of the correlation between IR and X-ray emission in our sample of optical TDEs can be estimated by considering a binomial calculation. The chance that a TDE has X-ray emission can be calculated as the detection fraction of the X-rays in our optical TDE sample (i.e. p=45/126=0.357). We could therefore test the null hypothesis (i.e. that each of the IR-flaring TDEs has an independent chance p=0.357 of X-ray detection) by computing the probability of seeing X-ray detections by chance:
Since , we reject the null hypothesis, meaning that the observed correlation of X-ray and IR emission in our IR-emitting optical TDE sample is likely not by chance. We can also convert the probability to sigma levels for a one-tailed normal distribution, which gives a 3.4 significance. We note that this analysis could be biased towards lower redshift sources, that are brighter and more easily observed by modern day telescopes.
@@ 6 Conclusions
We have compiled a catalogue of 134 confirmed TDEs up until the end of 2024. We collected multi-wavelength photometry (X-ray, UV, optical, and IR) along with publicly available spectra. The complete dataset is accessible via a dedicated GitHub repository and a Python application. Additionally, we created a list of strong TDE candidates.
@@
By implementing a custom Bayesian Blocks algorithm, we estimated the duration, rise time, decay time, and their ratio for the optical flares of 101 TDEs in our sample. A total of 103 flares were analysed, including all flares from repeating pTDEs (AT 2022dbl and AT 2020vdq). We find that the distributions of the ratio, durations, rise times, and decay times are well described by a log-normal distribution.
@@
The best-fit line for versus is given by , where and , with 95% confidence intervals of 0.768 to 1.062 and -0.65 to -0.04 for and respectively. This result indicates that the majority of the TDE population is confined within the scatter of this best-fit line. We also find a positive correlation between and . These results, in combination with the log-normal distributions of the timescales, can be used to create artificial light curves for future statistical studies.
@@
Our spectral analysis revealed that the majority of TDEs belong to the TDE-H+He class, followed by the TDE-H class, with the TDE-He and TDE-featureless occurring at slightly lower frequencies. This distribution is consistent with expectations based on the predominance of main-sequence stars in galactic centres. Although we computed timescale statistics for the different spectral classes, we did not identify any definite trends.
@@
We examined whether specific spectral classes tend to exhibit repeating flares, IR emission, or X-ray outbursts. Our analysis revealed that the spectral class distribution within these TDE sub-samples is unlikely to have arisen by random chance.
@@
We identified three new TDEs that show secondary flares in their optical light curves: AT 2024pvu, AT 2022exr, and AT 2021uvz. We further studied recently referenced TDEs displaying re-brightenings: AT 2020acka, and AT 2019ehz. We also compared the shape of the multiple flares for each TDE and found that their shapes are generally similar, excluding AT 2021mhg and AT 2020acka, which exhibit distinct differences.
@@
We observed a potential correlation between IR and X-ray emission. Most TDEs (14 out of 18) displaying IR flares also exhibit X-ray detections. We plan to explore this finding further in a future study. Moreover, all reported coronal line emitting TDEs were found to exhibit both X-ray and IR emission.
This comprehensive (mostly optical) TDE catalogue not only provides a robust dataset for statistical studies and machine learning applications but also paves the way for future population studies aimed at understanding the characteristics of still poorly understood TDEs, such as X-ray TDEs (Auchettl et al. 2017; Sazonov et al. 2021; Khorunzhev et al. 2022) and IR TDEs (Jiang et al. 2021; Panagiotou et al. 2023; Masterson et al. 2024). As we enter into the ‘data-rich era’ – especially with forthcoming surveys such as LSST, which are expected to capture tens of TDEs each night – this work serves as an essential tool for preparing for the challenges and opportunities ahead.
Data availability
Table 1 and the catalogue are available in electronic form at the CDS via anonymous ftp to http://cdsarc.u-strasbg.fr/ (130.79.128.5) or via https://cdsarc.cds.unistra.fr/viz-bin/qcat?J/A+A/. The data and python application (see Sect. 3), along with the aforementioned table information and overall catalogue, can be accessed through the TDECat GitHub repository202020https://github.com/dlangis/TDECat.
@@ Acknowledgements.
The authors thank the anonymous referee for comments and suggestions that helped improve this work. DAL, AP, and IL were funded by the European Union ERC-2022-STG - BOOTES - 101076343. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. KIIK has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002352, PI: M. Linares). NG acknowledge support by a grant from the Simons Foundation (00001470, NG). JD, DM, and ZT took part in this research under the auspices of the Science Internship Program at the University of California Santa Cruz. NG is grateful to Caitlyn Nojiri and Hannah Dykaar for their help. BHTOM has been based on the open-source TOM Toolkit by LCO and has been developed with funding from the European Union’s Horizon 2020 and Horizon Europe research and innovation programmes under grant agreements No 101004719 (OPTICON-RadioNet Pilot, ORP) and No 101131928 (ACME). We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts). This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogues from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. The CSS survey is funded by the National Aeronautics and Space Administration under Grant No. NNG05GF22G issued through the Science Mission Directorate Near-Earth Objects Observations Program. The CRTS survey is supported by the U.S. National Science Foundation under grants AST-0909182. This publication makes use of data products from NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. This work utilises Matplotlib (Hunter 2007), NumPy (Harris et al. 2020) and SciPy (Virtanen et al. 2020).References
- Discovery of an Outflow from Radio Observations of the Tidal Disruption Event ASASSN-14li. ApJ 819 (2), pp. L25. External Links: Document, 1510.01226, ADS entry Cited by: Appendix A.
- Radio detections of TDEs AT2020ksf and AT2020zso. Transient Name Server AstroNote 24, pp. 1–24. External Links: ADS entry Cited by: Appendix A.
- A very luminous jet from the disruption of a star by a massive black hole. Nature 612 (7940), pp. 430–434. External Links: Document, 2211.16530, ADS entry Cited by: Appendix A.
- A Continuum of H- to He-rich Tidal Disruption Candidates With a Preference for E+A Galaxies. ApJ 793 (1), pp. 38. External Links: Document, 1405.1415, ADS entry Cited by: Appendix A, §1, 6th item, 7th item.
- New Physical Insights about Tidal Disruption Events from a Comprehensive Observational Inventory at X-Ray Wavelengths. ApJ 838 (2), pp. 149. External Links: Document, 1611.02291, ADS entry Cited by: §6.
- Detection of an extremely soft X-ray outburst in the HII-like nucleus of NGC 5905.. A&A 309, pp. L35–L38. External Links: ADS entry Cited by: §1.
- The Zwicky Transient Facility: System Overview, Performance, and First Results. PASP 131 (995), pp. 018002. External Links: Document, 1902.01932, ADS entry Cited by: §1.
- The Broad Absorption Line Tidal Disruption Event iPTF15af: Optical and Ultraviolet Evolution. ApJ 873 (1), pp. 92. External Links: Document, 1809.07446, ADS entry Cited by: Appendix A, §1, 5th item, 9th item.
- iPTF16fnl: A Faint and Fast Tidal Disruption Event in an E+A Galaxy. ApJ 844 (1), pp. 46. External Links: Document, 1703.00965, ADS entry Cited by: §1.
- K-Corrections and Filter Transformations in the Ultraviolet, Optical, and Near-Infrared. AJ 133 (2), pp. 734–754. External Links: Document, astro-ph/0606170, ADS entry Cited by: §3.1.
- A Possible Relativistic Jetted Outburst from a Massive Black Hole Fed by a Tidally Disrupted Star. Science 333 (6039), pp. 203. External Links: Document, 1104.3257, ADS entry Cited by: Appendix A.
- ASAS-SN Discovery of A Probable Supernova in 2MASX J20390918-3045201. The Astronomer’s Telegram 7910, pp. 1. External Links: ADS entry Cited by: Appendix A.
- Swift J1112.2-8238: a candidate relativistic tidal disruption flare. MNRAS 452 (4), pp. 4297–4306. External Links: Document, 1507.03582, ADS entry Cited by: Table 1.
- Relativistic jet activity from the tidal disruption of a star by a massive black hole. Nature 476 (7361), pp. 421–424. External Links: Document, 1104.4787, ADS entry Cited by: Appendix A.
- The Swift X-Ray Telescope. Space Sci. Rev. 120 (3-4), pp. 165–195. External Links: Document, astro-ph/0508071, ADS entry Cited by: §1.
- The swift xrt data reduction guide. Version 1, pp. 28. Cited by: §B.1.1.
- An Ultraviolet Spectrum of the Tidal Disruption Flare ASASSN-14li. ApJ 818 (2), pp. L32. External Links: Document, 1601.03331, ADS entry Cited by: Appendix A.
- Swift J2058.4+0516: Discovery of a Possible Second Relativistic Tidal Disruption Flare?. ApJ 753 (1), pp. 77. External Links: Document, 1107.5307, ADS entry Cited by: Appendix A.
- The Pan-STARRS1 Surveys. arXiv e-prints, pp. arXiv:1612.05560. External Links: Document, 1612.05560, ADS entry Cited by: §1.
- A detailed spectroscopic study of tidal disruption events. A&A 659, pp. A34. External Links: Document, 2109.00016, ADS entry Cited by: Appendix A, Appendix A, Appendix A, Appendix A, Appendix A, Appendix A, Appendix A, Appendix A, §4.2.
- Spectroscopic observation of the rising, blue, featureless, nuclear transient AT2020vdq.. Transient Name Server AstroNote 115, pp. 1. External Links: ADS entry Cited by: footnote 16.
- The seven year Swift-XRT point source catalog (1SWXRT). A&A 551, pp. A142. External Links: Document, 1302.7113, ADS entry Cited by: §B.1.1.
- Long Fading Mid-infrared Emission in Transient Coronal Line Emitters: Dust Echo of a Tidal Disruption Flare. ApJ 832 (2), pp. 188. External Links: Document, 1605.05145, ADS entry Cited by: §1.
- First Results from the Catalina Real-Time Transient Survey. ApJ 696 (1), pp. 870–884. External Links: Document, 0809.1394, ADS entry Cited by: §1, §3.1.
- Smearing estimate: a nonparametric retransformation method. Journal of the American Statistical Association 78 (383), pp. 605–610. External Links: Document, Link, https://www.tandfonline.com/doi/pdf/10.1080/01621459.1983.10478017 Cited by: footnote 15.
- AT 2020nov: Evidence for Disk Reprocessing in a Rare Tidal Disruption Event. ApJ 983 (1), pp. 28. External Links: Document, 2412.12991, ADS entry Cited by: §1.
- Aproximations to the radii of Roche lobes.. ApJ 268, pp. 368–369. External Links: Document, ADS entry Cited by: §4.3.
- Candidate tidal disruption events from the XMM-Newton slew survey. A&A 462 (3), pp. L49–L52. External Links: Document, astro-ph/0612340, ADS entry Cited by: §1.
- An Analysis of the Shapes of Interstellar Extinction Curves. V. The IR-through-UV Curve Morphology. ApJ 663 (1), pp. 320–341. External Links: Document, 0705.0154, ADS entry Cited by: §3.3.
- The Pan-STARRS1 Database and Data Products. ApJS 251 (1), pp. 7. External Links: Document, 1612.05243, ADS entry Cited by: §1.
@@
- Sherpa: a mission-independent data analysis application. In Astronomical Data Analysis, J. Starck and F. D. Murtagh (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4477, pp. 76–87. External Links: Document, astro-ph/0108426, ADS entry Cited by: §B.4.
- Tidal Disruption Events Prefer Unusual Host Galaxies. ApJ 818 (1), pp. L21. External Links: Document, 1601.04705, ADS entry Cited by: §4.3.